# Quantum Rationale-Aware Graph Contrastive Learning for Jet Discrimination

### v6

@@ D.2.3 Overall QRGCL Objective
Appendix E Details of Benchmark Models
@@
We developed two classical models (GNN and EQGNN), three quantum models (QGNN, EQGNN, and QCL), five hybrid classical-quantum models (CQCL, 3 QCGCL variants, and QRGCL with $RX+H$ encoding) (see Table 3), and the classical counterpart of QRGCL, i.e., CRGCL (see Table 2).
E.1 Classical RGCL (CRGCL)
@@
The classical RG (CRG) of RGCL is a GNN estimator designed to generate node representations from graph-structured data, which acts as the counterpart of QRGCL. It consists of three graph convolutional layers (GCN) that reduce feature dimensionality from the input to 32, 16, and finally, 8. Each layer is followed by batch normalization to stabilize training and reduce internal covariate shifts. ReLU activation is applied after the first two layers, while dropout is used to mitigate overfitting. The final output passes through a linear layer to yield a single value per node, followed by a softmax for probabilistic representation. For a fair comparison between the CRG and QRG, we use the well-established ParticleNet as the encoder network, followed by the same projection head used in QRGCL.
E.2 GNN
The architecture consists of layers where the message-passing mechanism updates node features based on neighboring nodes and their relationships. An MLP processes the node features in each layer, while an edge MLP aggregates the edge attributes. After several layers, the updated node features are pooled to form a graph-level representation, passing through a fully connected neural network for classification. GNNs can inherently handle the permutation of nodes due to their graph-based nature.
E.3 Equivariant GNN (EGNN)
EGNNs extend GNNs by incorporating symmetry transformations, such as rotational and translational equivariance. This means that the model’s predictions remain consistent under transformations of the input data. The coordinates of the nodes are updated during each layer based on interactions with neighboring nodes. The model utilizes additional features like distance metrics between node pairs to enforce equivariance in graph processing.
E.4 Quantum GNN (QGNN)
@@
In QGNN, the graph’s node features are embedded into quantum states using an embedding layer. These quantum states evolve under a parameterized Hamiltonian, which encodes the interactions between nodes (qubits) based on the graph’s adjacency matrix. The QGNN model relies on unitary transformations to evolve the quantum state over multiple layers. After the final layer, quantum measurements are performed, and the results are passed through a fully connected layer to make predictions.
E.5 Equivariant Quantum GNN (EQGNN)
@@
EQGNNs [33, 3, 39, 40, 41] are quantum analogs of EGNNs, incorporating equivariance into the quantum architecture. Like QGNNs, EQGNNs operate on quantum states but ensure that the learned representations are equivariant under symmetries, such as permutations or rotations. The final quantum states are aggregated through pooling, ensuring the network remains permutation-equivariant. This aggregation is followed by classical post-processing to yield the model’s predictions.
E.6 QCL
@@
QCL employs a quantum convolutional neural network (QCNN) with input size $18\times 18$ with zero padding as the encoder, where data-reuploading circuits (DRCs) replace classical convolutional kernels. The QCNN consists of 2 similar layers: the first with a ($3\times 3$) kernel, stride 2, and a 3-qubit DRC as a filter. The final encoding size after flattening is 16. The projection network either uses a 2-node linear layer.
E.7 Classical-Quantum CL (CQCL)
Hybrid CQCL is similar to QCL except for the final projection head, which uses a $\log_{2}^{d=2}$ qubit (d = embedding dimension) DRC circuit. To evaluate the ability of QCL and CQCL to generate generalized representations, we make predictions using a simple MLP, an input layer of size 256, and a hidden layer of size 32, both with BatchNorm and LeakyReLU.
E.8 Quantum-Classical GCL (QCGCL)
@@
QCGCL architecture initially uses multiple layers of graph attention-based convolution (GATConv) layers to capture graph features. Each GAT layer is followed by batch normalization and residual connections, which allow the model to skip layers for improved training. The graph-level embedding is achieved through mean and max pooling. A quantum circuit then processes this pooled GNN output. We benchmarked QCGCL using 3 different quantum encoders, as depicted in Table 3. Finally, the GNN and quantum outputs are concatenated and passed through a fully connected readout layer to generate the final prediction.
Appendix F Details of Experimental Setup
F.1 Simulation Tools and Environment
We implemented all the models using the PyTorch 2.2.0 [27] framework for classical computations and Pennylane 0.38.0 [28] and TorchQuantum 0.1.8 [29] for quantum circuit simulation. We used the Deep Graph Library (DGL) 2.1.0+cu121 [31] to handle graph operations and the Qiskit 0.46.0 [30] framework to simulate quantum circuits. The computing infrastructure consisted of Intel(R) Xeon(R) CPUs (x86) with a clock frequency of 2 GHz, equipped with 4 vCPU cores and 30 GB of DDR4 RAM. For GPU acceleration, we leveraged two NVIDIA T4 GPUs, each with 2560 CUDA cores and 16 GB of VRAM, significantly boosting the performance of deep learning tasks.
F.2 Hyperparameters and Configurations
We varied the hyperparameters related to CRG, QRG, data augmentation, and training for the proposed QRGCL model, highlighting the test accuracy, AUC, and F1 scores across various encoder types, learning rates (LR), and entanglement strategies. We used Adam optimizer with a $1\times 10^{-3}$ learning rate across all the models and a binary cross-entropy loss function. 10-fold cross-validation was performed for each model, and the mean and standard deviation were calculated for each metric. The hidden feature size for classical GNN and EGNN was 10, while for the QGNN and EQGNN, it was $2^{n_{\alpha}}$=8. In the case of QCL and CQCL, $n_{layer}$ stands for the depth of the quantum circuit, and in other models, it refers to the number of GNN layers.
The choice of epochs and batch sizes varies based on the computational requirements of classical and quantum models. Classical models like GNN and EGNN use larger batch sizes (64) and fewer epochs (19) to leverage efficient gradient updates, allowing faster convergence. In contrast, quantum models such as QGNN and EQGNN use much smaller batch sizes (1) due to quantum hardware limitations, yet still require 19 epochs to ensure adequate learning despite more minor updates per step. For fully quantum and hybrid models like QRGCL and QCGCL, larger batch sizes (2000) and more epochs (50) are used due to the longer backpropagation caused by the complex custom loss function, with multiple loss components ensuring that the model has sufficient time to learn robust quantum-based representations, as the slower convergence can hinder effective learning.
Appendix G Details of Results and Discussion
We evaluated the performance of QRGCL against two classical models (GNN and EQGNN), three quantum models (QGNN, EQGNN, and QCL), and five hybrid classical-quantum models (CQCL, three variants of QCGCL, and QRGCL with RX+H encoding). The area under the curve (AUC) was selected as the benchmark metric due to its effectiveness in assessing binary classification performance across all thresholds. The number of trainable parameters of CRG was 1,073, compared to 45 for QRG.
The hyperparameters specific to QRGCL include encoder type, learning rate, and entanglement type, as detailed in Table 1. Among the encoders, $RX$ and $H$ achieved the highest AUC values at 76.30% and 76.71%, respectively, while displacement-amplitude and $RZ$ encodings performed poorly, with AUCs of 72.10% and 67.32%. The optimal learning rate of $1\times 10^{-3}$ produced the highest AUC of 76.20%, with higher rates resulting in decreased performance. The $SWAP$ entanglement type yielded the best overall results, achieving an AUC of 76.20%. $CNOT$ and $CZ$ entanglements performed strongly, while other configurations underperformed. For the proposed QGRCL, two possible combinations of $RX$ and $H$ encoders the learning rate of $1\times 10^{-3}$ and $SWAP$ entangler were tried.
For both the classical RGCL and QRGCL models, tunable parameters included the number of nodes per head ($n_{\alpha}$) ranging from 7 to 10, the number of classical or quantum layers ($n_{layer}$) ranging from 2 to 5, and the augmentation ratio ranging from 0.1 to 0.3. As shown in Table 2, QRGCL outperforms the classical RGCL with 8 and 10 nodes, achieving higher AUC values. At 3 layers, QRGCL surpasses the classical model, achieving 74.71% compared to 73.58%. However, it underperformed at 2 and 4 layers. With a 0.1 augmentation ratio, QRGCL significantly outperformed the classical approach, achieving 78.78% compared to 74.09%. Overall, the optimal parameters for QRGCL were found to be $n_{\alpha}=7$, $n_{layer}=3$, and an augmentation ratio of 10%.
Based on Table 3 and Figure 5b, it is evident that our proposed QRGCL model achieves the highest AUC score (77.53%). We took the best two well-performing encoders from Table 1 and developed two variants by hybridizing them, and further experimentations revealed that $H$ followed by $RX$ gate is more robust. Additionally, the relatively large number of parameters in the QRGCL is due to the utilization of the ParticleNet GNN encoder, which possesses 125k parameters, while the QRG circuit only has 45 parameters. Tables 1, 2, and 3 complement the ablation studies for QRGCL by presenting results for configurations without rationale-awareness (QGNN, GNN, EGNN, EQGNN), analyzing VQC components (such as encoding variants, entanglement structures, and variations in qubits and layers), exploring different classical-quantum interfaces (including quantum-only and hybrid architectures, as well as a classical-only baseline), and examining the effects of rationale-guided data augmentation. Figure 5a shows both training and testing losses decrease steadily in QRGCL, indicating practical training with a continuous reduction in error. By around 800 epochs, the losses stabilize, suggesting that the model has reached a convergence point where further training yields minimal improvements. The small gap between the two loss curves suggests minimal overfitting and generalization of test data. Both training and testing accuracies rise sharply within the first 200 epochs, reflecting rapid learning and an effective initial adjustment of model parameters. Around 800 epochs, both accuracies stabilize. The test accuracy levels off slightly above 70%, while training accuracy remains close but slightly lower. This stability indicates that QRGCL has reached its maximum performance capacity on this dataset, with limited fluctuations.

### v7

@@ B.2.3 Overall QRGCL Objective
B.3 Details of Quantum Feature Encoders
To evaluate the impact of different quantum feature maps on the performance of the QRG, we benchmarked several encoding strategies.
B.3.1 Angle Encoding (RX, RY, RZ, H)
Angle encoding embeds $N$ classical features $x=\{x_{1},\dots,x_{N}\}$ into the rotation angles of $N$ qubits, maintaining a parameter-efficient 1:1 correspondence between features and qubits. In the standard Rotation Encoding variants (RX, RY, RZ), the feature $x_{i}$ serves as the rotation parameter for a Pauli rotation gate acting on the $i$-th qubit. For instance, the $R_{X}$ encoding prepares the state $|\psi\rangle=\bigotimes_{i=1}^{N}R_{X}(x_{i})|0\rangle=\bigotimes_{i=1}^{N}\exp\left(-i\frac{x_{i}}{2}\hat{\sigma}_{x}\right)|0\rangle$. To increase the expressivity of this feature map, the H + Angle variant first places the qubits in a uniform superposition using Hadamard ($H$) gates before applying the rotations, resulting in the state $|\psi\rangle=\bigotimes_{i=1}^{N}R_{X}(x_{i})H|0\rangle$.
B.3.2 IQP Encoding
The Instantaneous Quantum Polynomial (IQP) encoding embeds features into a quantum state conjectured to be hard to simulate classically. This strategy diagonalizes the encoding in the $Z$-basis using Hadamard gates and specific interactions. The circuit structure consists of an initial layer of Hadamard gates, followed by a layer of diagonal phase gates that encode the features, and a final entangling layer. Mathematically, for a feature vector $x$, the encoding unitary is defined as $U_{IQP}(x)=\left(\prod_{(i,j)\in S}e^{ix_{i}x_{j}\hat{\sigma}_{z}^{(i)}\otimes\hat{\sigma}_{z}^{(j)}}\right)\left(\bigotimes_{i=1}^{N}e^{ix_{i}\hat{\sigma}_{z}^{(i)}}\right)\bigotimes_{i=1}^{N}H$, where $S$ represents the set of entangled qubit pairs. In our experiments, we utilized a standard repetition pattern ($n_{repeats}=1$) to embed the correlations between features.
B.3.3 Amplitude Encoding
Amplitude encoding is a space-efficient strategy that embeds a normalized classical vector $x$ of dimension $2^{N}$ into the probability amplitudes of an $N$-qubit entangled state. Given a normalized input vector such that $\|x\|^{2}=1$, the state is prepared as $|\psi\rangle=\sum_{i=0}^{2^{N}-1}x_{i}|i\rangle$, where $|i\rangle$ are the computational basis states. While this method theoretically allows for the encoding of exponentially many features into 7 qubits (up to $128$ features), it generally requires state preparation circuits that can scale exponentially in depth without specific optimization. We restricted the input to the graph node features for this study.
B.3.4 Displacement Encoding
[…]
Appendix C Details of Benchmark Models
@@
We developed two classical models (GNN and EGNN), three quantum models (QGNN, EQGNN, and QCL), five hybrid classical-quantum models (CQCL, 3 QCGCL variants, and QRGCL with $RX+H$ encoding) (see Table 1), and the classical counterpart of QRGCL, i.e., CRGCL (see Table 2). This section provides comprehensive architectural specifications and training configurations for each benchmark, ensuring reproducibility.
C.1 Classical RGCL (CRGCL)
C.1.1 Architecture
@@
The classical RG (CRG) of RGCL is a GNN estimator designed to generate node representations from graph-structured data, which acts as the counterpart of QRGCL. It consists of three graph attention network (GAT) layers with the following configuration: Layer 1 (input $\rightarrow$ 32 features, 4 attention heads), Layer 2 (32 $\rightarrow$ 16 features, 2 heads), and Layer 3 (16 $\rightarrow$ 8 features, 1 head). Each layer is followed by batch normalization (momentum = 0.1) to stabilize training and reduce internal covariate shifts. ReLU activation is applied after the first two layers, while dropout (p = 0.1) is used to mitigate overfitting. The final output passes through a linear layer to yield a single value per node, followed by a softmax for probabilistic node importance scoring. For a fair comparison between the CRG and QRG, we use the well-established ParticleNet as the encoder network, followed by the same projection head used in QRGCL.
C.1.2 Training Configuration
CRGCL employs a combined loss function: $\mathcal{L}_{\text{CRGCL}}=\mathcal{L}_{\text{RA}}+\lambda\mathcal{L}_{\text{CP}}+\alpha\mathcal{L}_{\text{align}}+\beta\mathcal{L}_{\text{uniform}}+\delta\mathcal{L}_{\text{InfoNCE}}$ with hyperparameters $\lambda=1.0$, $\alpha=0.5$, $\beta=0.0$ (uniformity term disabled), $\delta=0.5$, and temperature $T=0.1$ for all contrastive components. We use the Adam optimizer with learning rate $1\times 10^{-3}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$, with no weight decay. Training follows a two-stage procedure: (i) self-supervised pre-training for 50 epochs with batch size 2000 and gradient clipping (max norm 1.0), and (ii) supervised fine-tuning for 1000 epochs with frozen encoder parameters using BCE loss. Data augmentation includes node dropping (0.1), edge perturbation (0.05), and feature masking (0.1) with an augmentation ratio of 0.1.
C.2 GNN
C.2.1 Architecture
The baseline GNN architecture consists of 5 EdgeConv blocks with k=16 nearest neighbors per block. The architecture processes input features through: EdgeConv Block 1 (channels=[32,32,32]), EdgeConv Block 2 ([64,64,64]), EdgeConv Block 3 ([64,64,64]), EdgeConv Block 4 ([64,64,64]), and EdgeConv Block 5 ([10,10,10]). Each block contains a 3-layer MLP with BatchNorm and ReLU activation, followed by a max pooling operation over neighboring elements. The message-passing mechanism updates node features based on neighboring nodes using edge features constructed as $\Delta f_{ij}=f_{j}-f_{i}$. Residual connections are applied around each block. After all EdgeConv layers, global mean pooling aggregates node features into a graph-level representation, which is passed through a linear layer ([10 $\rightarrow$ 2]) with softmax for classification. GNNs can inherently handle node permutations due to their graph-based nature.
C.2.2 Training Configuration
[…]
C.3 Equivariant GNN (EGNN)
C.3.1 Architecture
EGNNs extend GNNs by incorporating SE(3) equivariance, ensuring predictions remain invariant to rotations and translations. The architecture consists of 4 EGNN layers with hidden dimensions [32, 64, 64, 10] and edge dimensions fixed at 32. Each layer updates both node features $h_{i}$ and coordinates $x_{i}\in\mathbb{R}^{3}$ through: (i) edge features $m_{ij}=\phi_{e}(h_{i}^{l},h_{j}^{l},\|x_{i}^{l}-x_{j}^{l}\|^{2},a_{ij})$, (ii) coordinate updates $x_{i}^{l+1}=x_{i}^{l}+\sum_{j}(x_{i}-x_{j})\cdot\phi_{x}(m_{ij})$, and (iii) node feature updates $h_{i}^{l+1}=\phi_{h}(h_{i}^{l},m_{i})$ where $m_{i}=\sum_{j}m_{ij}$. Each $\phi$ is a 2-layer MLP with SiLU activation and layer normalization. Initial coordinates are derived from kinematic features: $x=p_{T}\cos(\phi)$, $y=p_{T}\sin(\phi)$, $z=p_{T}\sinh(\eta)$. After all layers, global mean pooling over node features produces the graph representation, which is processed through an MLP readout ([10 $\rightarrow$ 32 $\rightarrow$ 2]).
C.3.2 Training Configuration
EGNN uses BCE loss with Adam optimizer (learning rate $1\times 10^{-3}$, weight decay $1\times 10^{-5}$). The learning rate follows CosineAnnealingLR schedule ($T_{\text{max}}$=19, $\eta_{\text{min}}=1\times 10^{-6}$) with 3-epoch linear warm-up from $1\times 10^{-5}$ to $1\times 10^{-3}$. Training runs for 19 epochs with a batch size of 64 and gradient clipping (max norm: 1.0). Layer normalization is applied instead of batch normalization to preserve equivariance properties. Coordinate clipping at $\pm 10.0$ prevents explosion. No dropout is used to maintain equivariance.
C.4 Quantum GNN (QGNN)
C.4.1 Architecture
@@
In QGNN, the graph’s node features are first processed through a classical MLP encoder ([8 $\rightarrow$ 16 $\rightarrow$ 8]) with ReLU and BatchNorm, followed by a linear projection ([8 $\rightarrow$ $n_{\alpha}$]) to map features to qubit count. These features are embedded into quantum states using a parameterized quantum circuit with $n_{\alpha}=3$ qubits and 6 variational layers. The quantum circuit structure per layer consists of: (i) Hadamard gates $H^{\otimes n_{\alpha}}$, (ii) RY rotation encoding with $\theta_{i}=\arctan(x_{i})$, (iii) parameterized U3 gates on each qubit with 3 parameters ($\alpha,\beta,\gamma$), and (iv) CNOT ladder entanglement with wrap-around connection. These quantum states evolve under the parameterized Hamiltonian, which encodes node interactions based on the graph’s adjacency matrix. The QGNN model utilizes unitary transformations to evolve the quantum state across multiple layers. After the final layer, quantum measurements are performed in the Z-basis, producing a $2^{n_{\alpha}}=8$ dimensional classical vector. The results are passed through an MLP classifier ([8 $\rightarrow$ 16 $\rightarrow$ 2]) with ReLU and Dropout (0.1) to make predictions. Total parameters: 5,156 (54 quantum + 5,102 classical).
C.4.2 Training Configuration
QGNN is trained using BCE loss with Adam optimizer (learning rate $1\times 10^{-3}$, no weight decay). Quantum circuit simulation uses PennyLane’s default qubit device with statevector simulation (no shots) and parameter-shift rule for differentiation. Due to quantum circuit overhead, the batch size is 1 with gradient accumulation over 64 steps (effective batch size = 64). Training runs for 19 epochs with gradient clipping (max norm 0.5) to prevent barren plateaus. Quantum parameters are initialized uniformly in $\pi/4,\pi/4]$, while classical parameters use Xavier initialization. Layer-wise learning rates are applied: first 2 quantum layers at $5\times 10^{-4}$, last 4 at $1\times 10^{-3}$, classical layers at $1\times 10^{-3}$. Gaussian parameter noise ($\sigma=0.01$) is added during training for robustness.
C.5 Equivariant Quantum GNN (EQGNN)
C.5.1 Architecture
@@
EQGNNs (Jahin_2025; forestano_comparison_2024; forestano_accelerated_2023; forestano_discovering_2023; forestano_identifying_2023) are quantum analogs of EGNNs, incorporating equivariance into the quantum architecture. The model begins with an equivariant classical encoder: Equivariant MLP Layer 1 ([8+3 $\rightarrow$ 16]) and Layer 2 ([16 $\rightarrow$ 16]), both with Layer Norm and SiLU activation. Node features $h_{i}\in\mathbb{R}^{8}$ and coordinates $x_{i}\in\mathbb{R}^{3}$ are jointly processed. Permutation-invariant pooling (DeepSets style) produces a fixed-size feature vector projected to $n_{\alpha}$ values. Like QGNNs, EQGNNs operate on quantum states (3 qubits, 6 layers, same circuit structure as QGNN) but ensure that the learned representations are equivariant under symmetries, such as permutations or rotations. The final quantum states are aggregated through pooling, ensuring the network remains permutation-equivariant. This aggregation is followed by classical post-processing to yield the model’s predictions. The model ensures: (i) permutation equivariance in classical encoding, (ii) graph isomorphism invariance in quantum embedding, and (iii) permutation invariance in final pooling. Total parameters: 5,140 (54 quantum + 5,086 classical).
C.5.2 Training Configuration
EQGNN uses BCE loss with Adam optimizer (learning rate $1\times 10^{-3}$, weight decay $1\times 10^{-5}$). The learning rate follows the StepLR schedule (step_size=5, gamma=0.5). Training runs for 19 epochs with batch size 1 and gradient accumulation over 64 steps. Gradient clipping is set to max norm 0.5. During training, periodic equivariance checks ensure $f(P\cdot G)\approx P\cdot f(G)$ for permutation $P$. Layer normalization is used instead of batch normalization to preserve equivariance, and no dropout is applied in equivariant layers. Coordinate noise augmentation with $\sigma=0.01$ is applied.
C.6 QCL
C.6.1 Architecture
@@
QCL employs a quantum convolutional neural network (QCNN) with input size $18\times 18$ (zero-padded from variable jet size) as the encoder, where data-reuploading circuits (DRCs) replace classical convolutional kernels. The QCNN consists of 2 quantum convolutional layers: Layer 1 with ($3\times 3$) kernel and stride 2, using a 3-qubit DRC as a filter, producing an $8\times 8\times 4$ feature map; Layer 2 with the same configuration, producing a $3\times 3\times 4$ feature map. Each DRC filter contains 3 qubits with 3 layers, where each layer includes: (i) feature encoding via RY($\theta_{\text{data}}$), (ii) trainable rotation RY($\theta_{\text{param}}$), and (iii) CNOT chain entanglement. Each filter has 18 parameters (3 qubits $\times$ 3 layers $\times$ 2 rotations). The final encoding size after flattening is 36 features, which are linearly projected to 16 dimensions. The projection network uses a 2-node linear layer ([16 $\rightarrow$ 2]). Total parameters: 280 (36 quantum + 244 classical).
C.6.2 Training Configuration
QCL uses InfoNCE contrastive loss: $\mathcal{L}_{\text{QCL}}=-\log[\exp(\text{sim}(z_{i},z_{j})/\tau)/\sum_{k}\exp(\text{sim}(z_{i},z_{k})/\tau)]$ where sim($\cdot$$\cdot$) is cosine similarity and $\tau=0.1$ (temperature). Positive pairs are the same jet with different augmentations, while negative pairs are different jets. Adam optimizer with learning rate $1\times 10^{-3}$ (no scheduler, no weight decay) is used. Training follows a two-stage procedure: pre-training for 1,000 epochs and fine-tuning for 500 epochs, with a batch size of 256 and 2 augmented views per sample. Data augmentation includes random crop ($16\times 16$ from $18\times 18$), random rotation ($\pm 15^{\circ}$), Gaussian noise ($\sigma=0.05$), and intensity scaling ($\pm 10\%$). Gradient clipping is set to a maximum norm of 1.0. Quantum circuits employ shot-based simulation (with 1024 shots) and parameter-shift differentiation. Quantum parameters are initialized uniformly in $\pi,\pi]$, with parameter sharing across filters for efficiency.
C.7 Classical-Quantum CL (CQCL)
C.7.1 Architecture
Hybrid CQCL uses a fully classical CNN encoder with a quantum projection head. The classical encoder consists of: Conv2D Layer 1 (1 $\rightarrow$ 16 channels, $3\times 3$ kernel, stride 2) with BatchNorm and ReLU, followed by Conv2D Layer 2 (16 $\rightarrow$ 32 channels, same configuration). After flattening ($32\times 4\times 4=512$ features), a linear layer projects to 256 dimensions with ReLU and Dropout (0.1). The quantum projection head utilizes amplitude encoding to embed the normalized 256-dimensional vector into 8 qubits ($\log_{2}(256)=8$). The variational circuit consists of three layers, each with RY and RZ rotations applied to all qubits, followed by linear CNOT entanglement (16 parameters per layer, totaling 48 parameters). Measurements are performed in the Z-basis on all qubits, producing an 8-dimensional output, which is finally projected to 2 dimensions. To evaluate the ability of QCL and CQCL to generate generalized representations, we make predictions using a simple MLP with an input layer of size 256 and a hidden layer of size 32, both with Batch Normalization and leaky ReLU. Total parameters: 250 (48 quantum + 202 classical).
C.7.2 Training Configuration
CQCL employs Supervised InfoNCE loss: $\mathcal{L}_{\text{CQCL}}=-\sum_{i}\log[\sum_{j\in\mathcal{P}(i)}\exp(\text{sim}(z_{i},z_{j})/\tau)/\sum_{k}\exp(\text{sim}(z_{i},z_{k})/\tau)]$ where $\mathcal{P}(i)$ includes all samples with the same label as $i$. Adam optimizer uses separate learning rates: $1\times 10^{-3}$ for classical layers and $5\times 10^{-4}$ for quantum layers, with weight decay $1\times 10^{-4}$. Training consists of 1000 pre-training epochs (batch size 256) and 500 fine-tuning epochs (batch size 128). Data augmentation includes all QCL augmentations plus Mixup ($\alpha=0.2$, classical features only) and CutOut ($4\times 4$ patches). Mixed-precision training is used: FP16 for classical, and FP32 for quantum. The quantum-classical interface requires normalization $z_{\text{norm}}=z/\|z\|$ for amplitude encoding, with hybrid optimization using alternating updates (5 classical steps: 1 quantum step).
C.8 Quantum-Classical GCL (QCGCL)
C.8.1 Architecture
@@
QCGCL architecture initially uses 6 GAT convolutional layers to capture graph features with the following configuration: GAT Layer 1 ([8 $\rightarrow$ 32], 4 heads, concat), Layer 2 ([128 $\rightarrow$ 64], 4 heads, concat), Layer 3 ([256 $\rightarrow$ 64], 2 heads, concat), Layer 4 ([128 $\rightarrow$ 32], 2 heads, concat), Layer 5 ([64 $\rightarrow$ 32], 1 head), and Layer 6 ([32 $\rightarrow$ 16], 1 head). Each GAT layer is followed by batch normalization, ReLU activation, and dropout (0.1). Residual connections are applied around each layer for improved training. The graph-level embedding is achieved through concatenated mean and max pooling, producing a 32-dimensional representation. A quantum circuit then processes this pooled GNN output. We benchmarked QCGCL using 3 different quantum encoders, as depicted in Table 1: (i) Variant 1 (RY+RX): 7 qubits with encoding $|0\rangle\rightarrow H\rightarrow RY(\theta_{i})\rightarrow RX(\phi_{i})$, 6 variational layers with RY($\alpha$) $\rightarrow$ RX($\beta$) and circular CNOT entanglement (84 parameters); (ii) Variant 2 (RY): same structure but only RY rotations (42 parameters); (iii) Variant 3 (Amplitude+RY): amplitude encoding followed by RY-based variational circuit (42 parameters). The classical embedding (32-d) and quantum embedding (7-d) are concatenated to form a 39-dimensional fused representation, which is processed through an MLP ([39 $\rightarrow$ 64 $\rightarrow$ 32 $\rightarrow$ 2]) with ReLU and Dropout (0.1), followed by softmax for final prediction. Total parameters: 7,448 (varies by quantum encoding).
C.8.2 Training Configuration
QCGCL uses a combined loss function: $\mathcal{L}_{\text{QCGCL}}=\lambda_{1}\mathcal{L}_{\text{InfoNCE}}+\lambda_{2}\mathcal{L}_{\text{BCE}}$ with $\lambda_{1}=0.5$ (contrastive weight) and $\lambda_{2}=0.5$ (classification weight), where InfoNCE uses temperature $\tau=0.1$. AdamW optimizer is employed with learning rate $1\times 10^{-3}$, weight decay $1\times 10^{-4}$, and CosineAnnealingWarmRestarts scheduler ($T_{0}$=10 epochs, $T_{\text{mult}}$=2, $\eta_{\text{min}}=1\times 10^{-6}$). Training runs for 50 epochs with a batch size of 128, a 5-epoch warm-up, and gradient clipping (max norm: 1.0). Data augmentation includes node dropping (0.1), edge perturbation (0.05), feature masking (0.1), and graph mixup ($\alpha=0.2$). A multi-view contrastive strategy uses 3 views (two classical augmented + one quantum-enhanced), with hard negative mining enabled (top 20% hardest negatives). Regularization includes dropout (0.1) in all layers, batch normalization after each GAT layer, attention dropout (0.1), and quantum parameter noise ($\sigma=0.01$). Quantum circuits utilize statevector simulation with parameter-shift differentiation and gate count optimization prior to execution.